Jan invested 1200 dollars to a certain simple intrest rate and 2200 dollars at a rate of 3 percent higher. Her annual earning was 253 dollars. Find the two intrest rates if she earned 121 dollars more on the higher than on the smaller.
Answer by checkley77(12844)

1200X+2200(X+3)=253 OR 1200X-253=-2200(X+3)
1200X+121=2200(X+3)NOW ADD THE 2 EQUATIONS
1200X-253=-2200(X+3)
---------------------
2400X-132=0
2400X=132
X=132/2400
X=.055 OR %5.5% FOR THE $1200 INVESTMENT.
.055+.03=.085 0R 8.5% FOR THE $2200 INVESTMENT.
PROOF
1200*.055+2200*.085=253
66+187=253
253=253